OpenStudy (asylum15):

The angular displacement θ radians of a flywheel at time t seconds is given by (10 marks) θ =9t2 −2t3. Find (i) the angular velocity and angular acceleration at time t and their values at t = 1 s (ii) the time when the angular acceleration is zero.

OpenStudy (anonymous):

(i)angular velocity at time t is 18t-6t2 n value at t=1 is 12 and angular acceleration at t is 18-12t and value at t=1 is 6 (ii) angular acceleration will be zero at t=3/2.